# Ordinary primes for $\operatorname{GL}_{2}$-type abelian varieties and weight $2$ modular forms

### v1

@@ 5.2. Density of $S(c)$
$D_{\ell}(g^{\prime},d,n,c)$ :=$:=\bigg{\{}(N_{i})_{1\leq i\leq g^{\prime}}\in\prod_{1\leq i\leq g^{\prime}}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{det}N_{i})^{n}=(\operatorname{det}N_{i^{\prime}})^{n}\text{ for all $1\leq i,i^{\prime}\leq g^{\prime}$}$
@@
Finally, the factor $n^{g^{\prime}d}$ in Equation 8 arises from the possible choices of $\zeta_{i,j}$. Hence, it suffices to estimate $|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ for $c\in\mu_{n}(\mathbb{F}_{\ell})$. Observe that if $g^{\prime}=1$, then by Lemma 5.5,
$|D_{\ell}(1,d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ =#$=\#\big{\{}N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\hskip 1.0pt}\overline{c_{0}}\cdot(\operatorname{det}N)^{d}\big{\}}$
$\leq\sum_{D\in\mathbb{F}_{\ell}^{\times}}\#\{N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\hskip 1.0pt}\overline{c_{0}}D^{d},\operatorname{det}N=D\}$
@@
For $g^{\prime}\geq 2$, we will compare the size of $D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})$ with the following matrix group
@@
$|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ $\leq\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{1}^{n}=\dots=\operatorname{det}N_{g^{\prime}}^{n},$
$\hskip 142.26378pt\prod_{1\leq i\leq g^{\prime}-1}(\operatorname{tr}N_{i})^{2d}\cdot(\operatorname{tr}N_{g^{\prime}})^{2d}=c{\hskip 1.0pt}\overline{c_{0}}\cdot(\operatorname{det}N_{g^{\prime}})^{d}\bigg{\}}$
$=\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{g^{\prime}}^{n}=D,\;T^{n}\cdot(\operatorname{tr}N_{g^{\prime}})^{2dn}=c{\hskip 1.0pt}\overline{c_{0}}D^{d},$
@@
Under the assumptions on the family of Galois representations in Theorem 1.6, the set $S(c)$ has density 00.
@@ 6.1. Applications to abelian varieties
By the assumptions, $A$ is a $\operatorname{GL}_{2}$-type abelian variety. By the proof of [Rib04, Theorem 2.1] (see also [Wu20, Proposition 1.5]), $A$ is isogenous over $\mathbb{Q}$ to $B^{r}$ for some simple $\operatorname{GL}_{2}$-type abelian variety $B$ over $\mathbb{Q}$. Moreover, by [Rib04, Theorem 2.1] and its proof, $K^{\prime}:=\operatorname{End}_{\mathbb{Q}}(B)\otimes\mathbb{Q}$ is a number field of degree equal to $\dim B$ with $K^{\prime}\hookrightarrow K\hookrightarrow\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$. Since $[K:\mathbb{Q}]=q$ (resp. $2q$), $[K^{\prime}:\mathbb{Q}]=1$ or $q$ (resp. $1$, $2$, $q$, or $2q$). If $[K^{\prime}:\mathbb{Q}]=1$, then $B/\mathbb{Q}$ is an elliptic curve. In particular, the set of ordinary primes of $B$ has positive density, and so does the set of ordinary primes of $A$.
Suppose that $[K^{\prime}:\mathbb{Q}]\neq 1$. By [Rib04, Theorem 4.4], under Serre’s modularity conjecture, $B$ is isogenous over $\mathbb{Q}$ to abelian varieties of the form $A_{f}$ associated to some newform $f$ of weight $2$ under the Eichler–Shimura theory. As Serre’s modularity conjecture is proved in [KW09a, KW09b], we can thus complete the proof by invoking Theorem 1.5, Proposition 3.12, and the isomorphism $K^{\prime}=\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}\simeq K_{f}$. ∎
@@
    for every primes $v\nmid N_{A}\ell$ of $F$, $\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))=X^{2}-(a_{v}(A)\otimes 1)X+(\epsilon(v)\mathrm{N}(v)\otimes 1)$, which is independent of $\ell$ (where $\epsilon:\operatorname{Gal}(\overline{F}/F)\rightarrow K^{\times}$ is some finite order character independent of $\ell$);
@@
Therefore, if we are in the case of (P2), then the result readily follows from Theorem 1.6 (with $T$ being the set of rational primes that split completely in $K$) and Proposition 3.8. In the case of (P1), we need to show that the set $\{v\in\Sigma_{F}:v\nmid N_{A}\text{ and }a_{v}(A)=0\}$ has density 00, which simply follows from Corollary 5.7 when $c=0$. The case of (P1) then follows from Theorem 1.7 and Proposition 3.8. ∎
@@ 6.2. Application to modular forms
    for every rational primes $p\nmid N\ell$, $\operatorname{det}(XI-\rho_{f,\ell}(\operatorname{Frob}_{p}))=X^{2}-(a_{p}(f)\otimes 1)X+(\epsilon(p)p\otimes 1)$, which is independent of $\ell$;
@@
    the set $\{p\nmid N:a_{p}(f)=0\}$ of rational primes has density 00;
@@
From these properties, we can already deduce the second part of Theorem 1.5 by applying Theorem 1.6 (with $T$ being the set of rational primes that split completely in $K_{f}$) or Theorem 1.7. The first part of Theorem 1.5, i.e., when $[K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$, now follows from Proposition 2.9 and the second part.
Appendix A Ordinary primes of CM modular forms of weight $2$
@@
$S_{0}:=\{v\in\Sigma_{F}:v^{\prime}|v\text{ for some degree $1$ prime $v^{\prime}\notin S^{\prime}$ of $F^{\prime}$}\}\subseteq S^{\mathrm{ord}},$
so it suffices to show that $S_{0}$ has density greater than or equal to $\frac{1}{[F^{\prime}:F]}$. Note that each $v\in S_{0}$ may split into at most $[F^{\prime}:F]$ primes of $F^{\prime}$ of degree $1$, so
@@
#$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{v^{\prime}\in\Sigma_{F^{\prime}}:\mathrm{N}(v^{\prime})\leq X\}.$
@@
First, since there exists an imaginary quadratic field $F^{\prime}$ such that $a_{p}(f)=0$ for all but finite many primes $p$ that are inert in $F^{\prime}$, it follows that the set of ordinary primes of $f$ has density at most $1/2$. For the lower bound, note that there exists an algebraic Hecke character $\psi$ of $F^{\prime}$ such that $\rho_{f,\lambda}\simeq\mathrm{Ind}_{F^{\prime}/\mathbb{Q}}\psi$ (after extending the scalar). In particular, there exists a number field $K$ (where the image of $\psi$ lies) such that for all but finite many primes $v^{\prime}$ of $F^{\prime}$, $\rho^{\prime}:=\rho_{f,\lambda}|_{\operatorname{Gal}(\overline{\mathbb{Q}}/F^{\prime})}$ is unramified at $v^{\prime}$ and $\operatorname{det}(XI-\rho^{\prime}(\operatorname{Frob}_{v^{\prime}}))$ splits into linear factor over $K$. As $f$ is of weight $2$, the data $(\rho_{f,\lambda},F^{\prime}/\mathbb{Q},K)$ satisfies the condition in Proposition A.1, so it follows that the set of ordinary primes of $f$ has density at least $1/2$. The result then follows. ∎

### v2

@@ 5.2. Density of $S(c)$
$D_{\ell}(g^{\prime},d,n,c)$ $:=\bigg\{(N_{i})_{1\leq i\leq g^{\prime}}\in\prod_{1\leq i\leq g^{\prime}}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{det}N_{i})^{n}=(\operatorname{det}N_{i^{\prime}})^{n}\text{ for all $1\leq i,i^{\prime}\leq g^{\prime}$}$
@@
Finally, the factor $n^{g^{\prime}d}$ in Equation 8 arises from the possible choices of $\zeta_{i,j}$. Hence, it suffices to estimate $|D_{\ell}(g^{\prime},d,n,c{\thinspace}\overline{c_{0}})|$ for $c\in\mu_{n}(\mathbb{F}_{\ell})$. Observe that if $g^{\prime}=1$, then by Lemma 5.5,
$|D_{\ell}(1,d,n,c{\thinspace}\overline{c_{0}})|$ $=\#\big\{N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\thinspace}\overline{c_{0}}\cdot(\operatorname{det}N)^{d}\big\}$
$\leq\sum_{D\in\mathbb{F}_{\ell}^{\times}}\#\{N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\thinspace}\overline{c_{0}}D^{d},\operatorname{det}N=D\}$
@@
For $g^{\prime}\geq 2$, we will compare the size of $D_{\ell}(g^{\prime},d,n,c{\thinspace}\overline{c_{0}})$ with the following matrix group
@@
$|D_{\ell}(g^{\prime},d,n,c{\thinspace}\overline{c_{0}})|$ $\leq\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg\{N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{1}^{n}=\dots=\operatorname{det}N_{g^{\prime}}^{n},$
$\hskip 142.26378pt\prod_{1\leq i\leq g^{\prime}-1}(\operatorname{tr}N_{i})^{2d}\cdot(\operatorname{tr}N_{g^{\prime}})^{2d}=c{\thinspace}\overline{c_{0}}\cdot(\operatorname{det}N_{g^{\prime}})^{d}\bigg\}$
$=\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg\{N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{g^{\prime}}^{n}=D,\;T^{n}\cdot(\operatorname{tr}N_{g^{\prime}})^{2dn}=c{\thinspace}\overline{c_{0}}D^{d},$
@@
Under the assumptions on the family of Galois representations in Theorem 1.7, the set $S(c)$ has density 0.
@@ 6.1. Applications to abelian varieties
By the assumptions, $A$ is a $\operatorname{GL}_{2}$-type abelian variety. By the proof of [Rib04, Theorem 2.1], $A$ is isogenous over $\mathbb{Q}$ to $B^{r}$ for some simple $\operatorname{GL}_{2}$-type abelian variety $B$ over $\mathbb{Q}$. Moreover, by [Rib04, Theorem 2.1] and its proof, $K^{\prime}:=\operatorname{End}_{\mathbb{Q}}(B)\otimes\mathbb{Q}$ is a number field of degree equal to $\dim B$ with $K^{\prime}\hookrightarrow K\hookrightarrow\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$. Since $K:\mathbb{Q}]=q$ (resp. $2q$), $K^{\prime}:\mathbb{Q}]=1$ or $q$ (resp. $1$, 22, $q$, or $2q$). If $K^{\prime}:\mathbb{Q}]=1$, then $B/\mathbb{Q}$ is an elliptic curve. In particular, the set of ordinary primes of $B$ has positive density, and so does the set of ordinary primes of $A$.
Suppose that $K^{\prime}:\mathbb{Q}]\neq 1$. By [Rib04, Theorem 4.4], under Serre’s modularity conjecture, $B$ is isogenous over $\mathbb{Q}$ to abelian varieties of the form $A_{f}$ associated to some newform $f$ of weight $2$ under the Eichler–Shimura theory. As Serre’s modularity conjecture is proved in [KW09a, KW09b], we can thus complete the proof by invoking Theorem 1.6, Proposition 3.12, and the isomorphism $K^{\prime}=\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}\simeq K_{f}$. ∎
@@
    for every prime $v\nmid N_{A}\ell$ of $F$, $\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))=X^{2}-(a_{v}(A)\otimes 1)X+(\epsilon(v)\mathrm{N}(v)\otimes 1)$, which is independent of $\ell$ (where $\epsilon:\operatorname{Gal}(\overline{F}/F)\rightarrow K^{\times}$ is some finite order character independent of $\ell$);
@@
Therefore, if we are in the case of (P2), then the result readily follows from Theorem 1.7 (with $T$ being the set of rational primes that split completely in $K$) and Proposition 3.8. In the case of (P1), we need to show that the set $\{v\in\Sigma_{F}:v\nmid N_{A}\text{ and }a_{v}(A)=0\}$ has density 0, which simply follows from Corollary 5.7 when $c=0$. The case of (P1) then follows from Theorem 1.8 and Proposition 3.8. ∎
@@ 6.2. Application to modular forms
    for every rational prime $p\nmid N\ell$, $\operatorname{det}(XI-\rho_{f,\ell}(\operatorname{Frob}_{p}))=X^{2}-(a_{p}(f)\otimes 1)X+(\epsilon(p)p\otimes 1)$, which is independent of $\ell$;
@@
    the set $\{p\nmid N:a_{p}(f)=0\}$ of rational primes has density 0;
@@
From these properties, we can already deduce the second part of Theorem 1.6 by applying Theorem 1.7 (with $T$ being the set of rational primes that split completely in $K_{f}$) or Theorem 1.8. The first part of Theorem 1.6, i.e., when $K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$, now follows from Proposition 2.9 and the second part.
Acknowledgments
We wish to thank Francesc Fité for answering our questions and providing helpful comments on Galois representations of $\mathrm{GL}_{2}$-type abelian varieties. We would like to thank Rachel Pries for the stimulating conversations. We also wish to thank the Max Planck Institute for Mathematics in Bonn for its support and inspiring atmosphere for collaborations. Lastly, we are grateful for the many helpful comments and suggestions the referees provided.
Data availability
There is no data used in this article.
Conflict of interest
The authors state that there is no conflict of interest.
Appendix A Ordinary primes of CM modular forms of weight 22
@@
[…]
so it suffices to show that $S_{0}$ has density greater than or equal to $\frac{1}{[F^{\prime}:F]}$. Note that each $v\in S_{0}$ may split into at most $F^{\prime}:F]$ primes of $F^{\prime}$ of degree $1$, so
@@
$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{v^{\prime}\in\Sigma_{F^{\prime}}:\mathrm{N}(v^{\prime})\leq X\}.$
@@
First, since there exists an imaginary quadratic field $F^{\prime}$ such that $a_{p}(f)=0$ for all but finite many primes $p$ that are inert in $F^{\prime}$, it follows that the set of ordinary primes of $f$ has density at most $1/2$. For the lower bound, note that there exists an algebraic Hecke character $\psi$ of $F^{\prime}$ such that $\rho_{f,\lambda}\simeq\mathrm{Ind}_{F^{\prime}/\mathbb{Q}}\psi$ (after extending scalars). In particular, there exists a number field $K$ (where the image of $\psi$ lies) such that for all but finite many primes $v^{\prime}$ of $F^{\prime}$, $\rho^{\prime}:=\rho_{f,\lambda}|_{\operatorname{Gal}(\overline{\mathbb{Q}}/F^{\prime})}$ is unramified at $v^{\prime}$ and $\operatorname{det}(XI-\rho^{\prime}(\operatorname{Frob}_{v^{\prime}}))$ splits into linear factors over $K$. As $f$ is of weight $2$, the data $(\rho_{f,\lambda},F^{\prime}/\mathbb{Q},K)$ satisfies the condition in Proposition A.1, so it follows that the set of ordinary primes of $f$ has density at least $1/2$. The result then follows. ∎